# Ultrafast Sampling-based Kinodynamic Planning via Differential Flatness

### 17 Mar 2026 v1

@@ VI-B Kinodynamic planning for high-d o f manipulators
Figure 8: Reactive planning with moving obstacles: our UR5 robot successfully performs a “pick, place, and reset” loop with our kinodynamic RRT-Connect planner without colliding with any of the static or moving obstacles. The obstacles are observed by an overhead Intel Realsense camera.
@@
VI-C Real experiments
Finally, we verify the benefit of our approach for online motion planning on a real Universal Robots (UR5) platform in a cluttered environment. The environment consists of multiple static and dynamic objects, perceived by an Intel Realsense camera with a top-down view. The depth image from the camera is used to generate a set of spheres stored in a collision-affording point tree (CAPT) compatible with simd-parallelized motion planning [ramsey2024-capt]. The spheres represent the obstacles in the environment in our Alg. 3. Our kinodynamic RRT-Connect planner in Sec. VI-A is used to find a trajectory from the start configuration to the goal.
VI-C1 Kinodynamic Planning versus Geometric Planning
We first consider a “pick and place” scenario in Fig. 1, where the robot generates a motion plan from its start position to a “pick” position near a granola box, and then move to a “place” position above a basket. The environment is cluttered with multiple objects, creating narrow passages such that a minor deviation from the planned trajectory can cause collisions. Therefore, accurate tracking is crucial for the safe realization of this “pick and place” task. The baseline is a geometric path, generated by a SIMD-parallelized RRT-Connect planner from VAMP [thomason2024vamp]. The path is sampled with a constant open-loop velocity, chosen such that the task’s duration is similar to ours for a fair comparison. A closed-loop velocity control input is computed by Ruckig [berscheid2021ruckig] and sent to the UR5 robot to execute.
Fig. 6 plots the tracking error during the experiments. Our trajectory is dynamically feasible, allowing the robot to smoothly and accurately execute the task without any collisions (Figs. 6(a) and 7(a)). Meanwhile, the geometric path has sharp turns, leading to overshoots and fluctuations (Fig. 6(b)). As a result, the robot collides with two nearby boxes while trying to turn, as seen in Fig. 7(b). The overshoots can be observed in Fig. 6(c) with two spikes in total tracking error $\|\mathbf{q}_{actual}-\mathbf{q}_{desired}\|$, at times $t\approx 4$s and $t\approx 9$s, where $\mathbf{q}_{actual}$ and $\mathbf{q}_{desired}$ denote the actual and desired values of the joint angles from the motion plans. Meanwhile, our tracking error stays low without any spikes, highlighting the benefits of having a dynamically feasible trajectory from our kinodynamic planner for accurate task executions.
VI-C2 Reactive Kinodynamic Planning
@@
VII Conclusions
This paper develops AkinoPDF, an ultrafast sampling-based kinodynamic planning technique, by solving the two-point boundary value problem and dynamics propagation in closed forms for a common class of differentially flat robot platforms, and performing extremely fast forward kinematics and collision checking via simd parallelism, available on most consumer CPUs. Our approach is exact, general, and able to generate a dynamically feasible trajectory in the range of microseconds to milliseconds, which is suitable for online and reactive planning in dynamic environments. Our method outperforms common sampling-based and optimization-based kinodynamic planners as well as time-parameterization of a geometric path in terms of planning times, offering a fast, reliable, and feasible solution for trajectory generation. We believe our approach unfolds a promising and exciting research area where it is possible to transform existing sampling-based motion planners into fast kinodynamic versions while only requiring general-purpose and widely available CPUs. This ability potentially can lead to a wide application of our method on different robots, in different settings, and for various tasks. It also spawns a challenging but intriguing research question of analyzing the optimality guarantees of our approach in the original state space, e.g., by combining the property of differential flatness with existing guarantees [karaman2011optimalmp] or by combining with an optimization-based planner. A limitation of our approach is that the robot system has to be differentially flat, which raises an interesting question of how we can extend our method to non-flat systems.
VIII Appendices
VIII-A Derivation of closed-form solutions in Example 4

### 21 Apr 2026 v2

@@ VII-C1 Kinodynamic Planning versus Geometric Planning
Figure 9: Reactive planning with moving obstacles: our UR5 robot successfully performs a “pick, place, and reset” loop with our FLASK-RRTConnect planner without colliding with any of the static or moving obstacles. The obstacles are observed by an overhead Intel Realsense camera.
@@
VII-C2 Reactive Kinodynamic Planning
@@
VIII Conclusions
This paper develops FLASK, an ultrafast sampling-based kinodynamic planning framework, by solving the two-point boundary value problem and dynamics propagation in closed forms for a common class of differentially flat robot platforms, and performing extremely fast forward kinematics and collision checking via simd parallelism, available on most consumer CPUs. Our approach is exact, general and can be applied to any sampling-based planners while offering theoretical guarantees on probabilistic exhaustivity and asymptotic optimality. It is able to generate a dynamically feasible trajectory in the range of microseconds to milliseconds, which is suitable for online and reactive planning in dynamic environments. Our method outperforms common sampling-based and optimization-based kinodynamic planners as well as time-parameterization of a geometric path in terms of planning times, offering a fast, reliable, and feasible solution for trajectory generation.
Our approach unfolds a promising and exciting research area where it is possible to transform existing sampling-based motion planners with theoretical guarantees into fast kinodynamic versions while only requiring general-purpose and widely available CPUs. As we have closed-form BVP solutions for a large class of nonlinear differentially flat systems, we can largely bypass the propagation-based planners and can even enable kinodynamic planning in challenging settings, e.g., with graph-based planners. This ability potentially can lead to a wide application of our method on different robots, in different settings, and for various tasks. Our method can quickly bootstrap optimization-based planners with quality and dynamically feasible initial solutions, especially when their objective function differs from our cost. Moreover, the ability to specify a duration for our trajectory offers an exciting potential integration task and motion planning with temporal constraints such as task deadlines.
IX Acknowledgement
We would like to thank J. Arden Knoll, Emiliano Flores, and Yitian Gao for their help with the baselines.
X Appendices
X-A Derivation of closed-form solutions in Example 4